OpenStudy (anonymous):

(1)/(x+2) + (x-2)/(x-5) + (4x+3)/(x^2-3x-10)

OpenStudy (shamim):

ok. i m trying to solve it

OpenStudy (shamim):

\[\frac{ 1 }{ x+2 }+\frac{ x-2 }{ x-5 }+\frac{ 4x+3 }{ x ^{2}-3x-10 }\]

OpenStudy (shamim):

\[=\frac{ 1 }{ x+2 }+\frac{ x-2 }{ x-5 }+\frac{ 4x+3 }{ x ^{2}-5x+2x-10 }\]

OpenStudy (shamim):

r u following me

OpenStudy (shamim):

\[=\frac{ 1 }{ x+2 }+\frac{ x-2 }{ x-5 }+\frac{ 4x+3 }{ (x+2)(x-5) }\]

OpenStudy (shamim):

\[=\frac{ x-5+x ^{2}-4+4x+3 }{ (x+2)(x-5) }\]

OpenStudy (shamim):

and so on

OpenStudy (shamim):

if u need more help plz ask me